OpenStudy (anonymous):

What is the x-intercept of the line with the equation -2x+1/2y=18 please help ive been stuck @myininaya @jim_thompson5910 @ganeshie8 @zepdrix @.Sam. @dan815 @abb0t

ganeshie8 (ganeshie8):

do you know that the given equation represents graph of a line ?

OpenStudy (anonymous):

no please help me

ganeshie8 (ganeshie8):

Okay, for x-intercepts you can do this: plugin y = 0 in the equation and solve x

ganeshie8 (ganeshie8):

-2x+1/2y=18 plugin y = 0 -2x + 1/2(0) = 18 -2x = 18 x = ?

OpenStudy (anonymous):

-9?

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

\(\large \color{red}{\checkmark}\)

OpenStudy (anonymous):

thanks can u help me with 2 more? @ganeshie8